Im stuck on this one :/
A rectangle is 5 feet longer than it is wide. If its perimeter is 22 feet, what is the dimension of the rectangle?
X= width
Length= X+5
Perimmeter formula is P=2L+2W

Answer by rfer(16322)

2x+2(x+5)=22
2x+2x+10=22
4x=12
x=3
x+5=8